# and the homomorphism conjecture for $\mathcal{A}_{g}$

### v1

@@ 5.4. The tautological ring of $\mathcal{X}_{g}^{s}$
The vanishing $\theta^{g+1}=0\in\mathsf{CH}^{g+1}(\mathcal{X}_{g})$ holds under our normalization conventions as explained in [34, Theorem 2.1], see also [16, 35]. For $s=2,$ the vertical tautological ring was studied and shown to be Gorenstein in [34, Theorem 3.3]. The tautological ring for families of abelian varieties was also considered in [3, Section 9.4] and [32, Section 6], and $\mathsf{R}^{*}(\mathcal{X}^{s}_{g})$ coincides with the definition of [3] for the universal family.
@@
    $X^{s}=X\times_{S}\times\ldots\times_{S}X$ is the $s$-fold fibered product, and $f^{s}:X^{s}\to S$ is the canonical morphism,
@@ 5.6. Proof of Theorem 13
$\underbrace{(\mathcal{A}_{r}\times\mathcal{X}_{g-r})\times_{\mathcal{A}_{g}}\ldots\times_{\mathcal{A}_{g}}(\mathcal{A}_{r}\times\mathcal{X}_{g-r})}_{s}\to\mathcal{X}_{g}^{s}$
@@ 6.2. Abel-Jacobi maps
where $\overline{\mathcal{A}}_{g+s}$ is a toroidal compactification for which $\mathsf{Tor}$ extends and $\mathcal{T}$ is a torus-bundle over $\mathcal{X}_{g}^{s}$. The composite $\mathcal{M}^{\mathrm{ct}}_{g,2s}\rightarrow\mathcal{T}\rightarrow\mathcal{X}_{g}^{s}$ is $\operatorname{aj}_{\mathsf{B}}$. A proof of the above factorization for $s=1$ is given in [41, Lemma 5.1], but the argument applies to any $s$.
@@ 6.3. The Abel-Jacobi pullback of $\operatorname{PR}_{g,s}$
where $\mathcal{M}_{1}^{\operatorname{ct,un}}(\pi_{1},(1),\ldots,(1))]^{\mathrm{vir}}$ equals the fundamental class $\mathcal{M}_{1,m}^{\mathrm{ct}}]$, and $I_{g_{i},N_{i}}(z)$ is the marked $I$-function, see [49, Proposition 1] for the exact expression in terms of tautological classes. As proven in Section 6.2, the Abel-Jacobi pullbacks preserve tautological classes
@@ 6.4. The Gorenstein kernels and generalized products
    $\Delta_{g,s}\neq 0$ for $(g,s)\in\{{(6,0)},(5,1)\}\,.$
By [8, Theorem 7], $\mathsf{K}_{5,2}$ is 1-dimensional. Therefore, Theorem 11 is a consequence of parts (i) and (ii) of Theorem 47 for $\Delta_{5,1}\in\mathsf{K}_{5,2}$. The only obstruction to cover all the cases of (66) is computational. At the moment, the remaining cases require too much time and memory to compute.
As a consequence of the non-vanishings in part (ii) of Theorem 47, we obtain the following result.
@@
The class $\operatorname{PR}_{5,1}]\in\mathsf{CH}^{5}(\mathcal{X}_{5})$ is not tautological.
@@
We now turn to the proof of Theorem 47 (i). The verification of (ii) is through $\mathsf{admcycles}$, and the corresponding code can be found in [27].
@@
where the last equality is a special case of [53, Proposition 6.8].

### v2

@@ 5.5. The tautological ring of $\mathcal{X}_{g}^{s}$
The proofs of Propositions 36 and 37 carry over to families of abelian varieties, hence the same pushforward and pullback properties also hold for the tautological rings $\mathsf{R}^{*}(\mathcal{X}_{g}^{s})$ and $\mathsf{R}^{*}_{\mathsf{vert}}(\mathcal{X}_{g}^{s})$.
@@
The vanishing $\theta^{g+1}=0\in\mathsf{CH}^{g+1}(\mathcal{X}_{g})$ holds under our normalization conventions as explained in [34, Theorem 2.1], see also [16, 35]. For $s=2,$ the vertical tautological ring was studied and shown to be Gorenstein in [34, Theorem 3.3]. The tautological ring for families of abelian varieties was also considered in [3, Section 9.4], [32, Section 6] and [48]. The tautological ring $\mathsf{R}^{*}(\mathcal{X}^{s}_{g})$ coincides with the definition of [3] for the universal family.
@@
    $X^{s}=X\times_{S}\cdots\times_{S}X$ is the $s$-fold fibered product, and $f^{s}:X^{s}\to S$ is the canonical morphism,
@@ 5.7. Proof of Theorem 12
$\underbrace{(\mathcal{A}_{r}\times\mathcal{X}_{g-r})\times_{\mathcal{A}_{g}}\cdots\times_{\mathcal{A}_{g}}(\mathcal{A}_{r}\times\mathcal{X}_{g-r})}_{s}\to\mathcal{X}_{g}^{s}$
@@ 6.2. Abel-Jacobi maps
where $\overline{\mathcal{A}}_{g+s}$ is a toroidal compactification for which $\mathsf{Tor}$ extends and $\mathcal{T}$ is a torus bundle over $\mathcal{X}_{g}^{s}$. The composite $\mathcal{M}^{\mathrm{ct}}_{g,2s}\rightarrow\mathcal{T}\rightarrow\mathcal{X}_{g}^{s}$ is $\operatorname{aj}_{\mathsf{B}}$. A proof of the above factorization for $s=1$ is given in [41, Lemma 5.1], but the argument applies to any $s$.
@@ 6.3. The Abel-Jacobi pullback of $\operatorname{PR}_{g,s}$
where $\mathcal{M}_{1}^{\operatorname{ct,un}}(\pi_{1},(1),\ldots,(1))]^{\mathrm{vir}}$ equals the fundamental class $\mathcal{M}_{1,m}^{\mathrm{ct}}]$, and $I_{g_{i},N_{i}}(z)$ is the marked $I$-function, see [50, Proposition 1] for the exact expression in terms of tautological classes. As proven in Section 6.2, the Abel-Jacobi pullbacks preserve tautological classes
@@ 6.4. The Gorenstein kernels and generalized products
    $\Delta_{g,s}\in\mathsf{K}_{g,2s}$ for all $g\leq 4$ and $s=2$,
  • (iii)
    $\Delta_{g,s}\neq 0$ for $(g,s)\in\{{(6,0)},(5,1),(4,2)\}\,.$
By [8, Theorem 7], $\mathsf{K}_{5,2}$ is 1-dimensional. Therefore, Theorem 10 is a consequence of parts (i) and (iii) of Theorem 48 for $\Delta_{5,1}\in\mathsf{K}_{5,2}$. Theorem 48 covers the first three cases listed in (70). The only obstruction to covering all the cases of (70) is computational. At the moment, the remaining cases require too much time and memory to compute.
As a consequence of the non-vanishings in part (iii) of Theorem 48, we obtain the following results.
@@
The classes $\operatorname{PR}_{5,1}]\in\mathsf{CH}^{5}(\mathcal{X}_{5})$ and $\operatorname{PR}_{4,2}]\in\mathsf{CH}^{5}(\mathcal{X}^{2}_{4})$ are not tautological.
@@
Calculations suggest that Speculation 50 is correct at least for $\mathsf{K}_{g\leq 7,0}$. Pixton suggests that the first interesting case of Speculation 50 is for codimension 7 cycles in $\mathsf{K}_{8,0}$.
6.5. Proof of Theorem 48
The verifications of Thereom 48 (ii) and (iii) are through $\mathsf{admcycles}$, and the corresponding code can be found in [27]. In these cases, we know the 3-spin relations are complete [8, Theorem 7, proof of Proposition 33]. With the current computational tools, the proof that $\Delta_{4,2}\neq 0$ requires additional optimization to produce the matrix of 3-spin relations, which is explained in [8, Section 7] and was used for [8, proof of Proposition 33].
@@ 6.5. Proof of Theorem 48
where the last equality is a special case of [54, Proposition 6.8].